OpenStudy (anonymous):

analytically show that the functions are inverse functions. f(x)=e^2x g(x)=ln sqr rt(x)

OpenStudy (asnaseer):

\[y=e^{2x}\]take logs of both sides to get:\[\ln(y)=2x\]rearrange to get:\[x=\frac{1}{2}\ln(y)=\ln(y^\frac{1}{2})=\ln(\sqrt{y})\]

OpenStudy (asnaseer):

therefore:\[g(x)=f^{-1}(x)\]

OpenStudy (asnaseer):

do you follow?

OpenStudy (anonymous):

not exactly. i mean i understand your math, i'm just confused on the whole inverse thing...

OpenStudy (asnaseer):

if f(x) is some function which returns a value for each value of 'x' you plug into it. the inverse would mean, if you are given the value of f(x), then what value of x was used to get that.

OpenStudy (asnaseer):

so in the first part of the proof, I got:\[x=\ln(\sqrt{y})=\ln(\sqrt{f(x)})\] what this is saying is that given a value for f(x), work out what x should be to get that value.

OpenStudy (anonymous):

so is \[f ^{-1}(x)=g(x)\] and \[g ^{-1}(x)=f(x)\]

OpenStudy (asnaseer):

yes